OpenStudy (anonymous):

simplify the following : a. csc teta sin teta ___________ b. cot teta / cos teta ____________ c.sec teta / tan teta _____________ d. cot teta / tan teta ____________

OpenStudy (pratyush5):

a. = 1 b.= 1/sin c.= 1/sin d.=cos^2/sin^2